# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ VII.1 What is yet to be nailed down theoretically
Experiments on an ensemble of small spin-glass grains, where $\xi_{\mathrm{micro}}$ would be of the order of the grains’ size, open up another avenue. Under these conditions, large sample-to-sample fluctuations may unveil exotic phenomena such as multifractality [32]. Reliably measuring single-grain magnetizations would be a (not minor) technical challenge but even the collective magnetization of a (small) set of grains may be useful to investigate finite-size effects experimentally. This will bring us closer to the broader field of structural glasses, where non-linear responses unveil length scales akin to $\xi_{\mathrm{micro}}$ [1, 7].
@@ VII.2 Experimental opportunities and challenges
Rejuvenation is nearly always attributed to temperature chaos. Is this true? An alternate view has been proposed by Berthier and Bouchaud [53, 54], who “demonstrated in a numerical simulation of the 4D Ising spin glass that a temperature change $\Delta T/T_{\mathrm{g}}$ induces strong rejuvenation effects, whereas the direct observation of the overlap between configurations reveals no sign of chaos on the dynamically relevant length scales.” Making reference to the claim of Jönsson et al. [163] that temperature chaos is the origin of rejuvenation, they state that this analysis “cannot be viewed as definitive evidence for temperature chaos.” Note that in Sec. 41 we point out that Jönsson et al. [163] obtained a value $1/(D_{\text{s}}/2-y)\approx 2.6$ whereas Berthier and Bouchaud [53] use $\/(D_{\text{s}}/2-y)\approx 1$, the value found in most simulations and by Zhai et al. [301]. This issue may be illuminated by the experiments suggested above.
@@
With the development of high magnetic field facilities such as the High Magnetic Field Laboratory at Florida State University, it is possible to fully saturate a dilute spin glass that might give insight to the interactions between spins. There has been a preliminary unpublished effort on CuMn to explore such. In the simplest case, deviations from a Brillouin function could be modeled to determine the interactions. In the preliminary study, it would appear such an analysis could provide information on at least nearest-neighbor interactions. This work is in its infancy but the initial results look promising.
@@ Acknowledgments
This work was partially supported by Ministerio de Ciencia, Innovación y Universidades (MICIU, Spain), Agencia Estatal de Investigación (AEI, Spain), and European Regional Development Fund (ERDF, A way of making Europe) through Grants no. PID2020-112936GB-I00 and PID2022-136374NB-C21. EM acknowledges the hospitality of Mark Bowick at UCSB (USA), where part of this work was done, and the support by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). His research was supported by the 2021 FIS (Fondo Italiano per la Scienza) funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) and the 2022 PRIN funding scheme (2022LMHTET - Complexity, disorder and fluctuations: spin glass physics and beyond) from Italian MUR (Ministry of University and Research). EDD and RLO were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Science and Engineering Division, under Award No. DE-SC0013599. The single CuMn crystals were grown by Dr. D.L. Schlagel at Ames Laboratory, which is operated by the U.S. DOE by Iowa State University under Contract No. DE-AC02-07CH11358. IGAP and FRT acknowledge the support of ICSC - Italian Research Center on High-Performance Computing, Big Data, and Quantum Computing, funded by the European Union - NextGenerationEU.

### v2

@@ VII.1.4 Experiments with potential for strong impact on theory
Experiments on an ensemble of small spin-glass grains, where $\xi_{\mathrm{micro}}$ would be of the order of the grains’ size, open up another avenue. Under these conditions, large sample-to-sample fluctuations may unveil exotic phenomena such as multifractality [34]. Reliably measuring single-grain magnetizations would be a (not minor) technical challenge but even the collective magnetization of a (small) set of grains may be useful to investigate finite-size effects experimentally. This will bring us closer to the broader field of structural glasses, where nonlinear responses unveil length scales akin to $\xi_{\mathrm{micro}}$ [1, 7, see also Sec. II.7.1].
@@ VII.2 Experimental opportunities and challenges
Rejuvenation is nearly always attributed to temperature chaos. Is this true? An alternate view has been proposed by Berthier and Bouchaud [56, 57], who “demonstrated in a numerical simulation of the 4D Ising spin glass that a temperature change $\Delta T/T_{\mathrm{g}}$ induces strong rejuvenation effects, whereas the direct observation of the overlap between configurations reveals no sign of chaos on the dynamically relevant length scales.” Making reference to the claim of Jönsson et al. [192] that temperature chaos is the origin of rejuvenation, they state that this analysis “cannot be viewed as definitive evidence for temperature chaos.” Note that in Sec. 42 we point out that Jönsson et al. [192] obtained a value $1/(D_{\text{s}}/2-y)\approx 2.6$ whereas Berthier and Bouchaud [56] use $(D_{\text{s}}/2-y)\approx 1$, the value found in most simulations and by Zhai et al. [375]. This issue may be illuminated by the experiments suggested above.
@@
Figure 56: Magnetization against the applied field for five CuMn alloys. The Mn concentrations are 2.6% (violet), 4% (gold), 7% (red), 7.92% (green) and 13.5% (black). Each sample has several temperatures plotted with temperatures ranging from at least 5 K above $T_{\mathrm{g}}$ to about 10 K. For example, the 13.5% sample ($T_{\mathrm{g}}=58$ K) temperature range was from 70 K to 4.3 K and the 2.6% sample ($T_{\mathrm{g}}=19$ K) temperature range was from 25 K to 4.3 K. Data from Dahlberg et al. [105].
With the development of high magnetic field facilities such as the High Magnetic Field Laboratory at Florida State University, it is possible to fully saturate a dilute spin glass that could give insight into the interactions between spins. There has been a preliminary unpublished effort on CuMn to explore such. The surprise is there is minimal temperature dependence in the $M$ vs $H$ data for the five CuMn samples investigated with concentrations varying from 2.6% Mn to 13.5% Mn (see Fig. 56 and previous related work in Fig. 7). This includes temperatures from at least 5 K above $T_{\mathrm{g}}$ to the order of 10 K. This work is in its infancy with an extensions of measurements on much lower Mn concentrations in progress.
@@ Acknowledgments
The authors thank Jean-Philippe Bouchaud and the anonymous referees for their helpful suggestions. This work was partially supported by Ministerio de Ciencia, Innovación y Universidades (MICIU, Spain), Agencia Estatal de Investigación (AEI, Spain, MCIN/AEI/10.13039/501100011033), and European Regional Development Fund (ERDF, A way of making Europe) through Grants no. PID2020-112936GB-I00 and PID2022-136374NB-C21. EM acknowledges the hospitality of Mark Bowick at UCSB (USA), where part of this work was done, and the support by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). His research was supported by the 2021 FIS (Fondo Italiano per la Scienza) funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) and the 2022 PRIN funding scheme (2022LMHTET - Complexity, disorder and fluctuations: spin glass physics and beyond) from Italian MUR (Ministry of University and Research). EDD and RLO were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Science and Engineering Division, under Award No. DE-SC0013599. The single CuMn crystals were grown by Dr. D.L. Schlagel at Ames Laboratory, which is operated by the U.S. DOE by Iowa State University under Contract No. DE-AC02-07CH11358. IGAP and FRT acknowledge the support of ICSC - Italian Research Center on High-Performance Computing, Big Data, and Quantum Computing, funded by the European Union - NextGenerationEU.